Circular No. 5732

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

α PISCIS AUSTRINI S. A. Stern, Southwest Research Institute, San Antonio; M. C. Fes-

tou, Observatoire Midi-Pyrenées, Toulouse; and D. A. Weintraub, Vanderbilt University, Nashville, report: “Mapping observations on Feb. 20-21 revealed 1.3-mm continuum emission in a broad, disk-like region around the nearby, main-sequence star α PsA (Fomalhaut, A3V; distance 6.7 pc). The observations supporting this discovery were made using the 7-channel MPIfR bolometer of the IRAM 30-m telescope on Pico Vileta, with a halfpower bandwidth at 1.3 mm of 12′′. The emission geometry appears to be a tilted disk with the position angle of the major axis near 100^{o} and an aspect ratio of 2:1; the major-axis emission exceeds 20 mJy at the 450-500-AU contour (∼ 1′ from the star). Additional emission may be present at larger distances. The peak 1.3-mm emission detected is 50 mJy, centered on the line of sight to the star. The emission is ascribed to an assemblage of cold, orbiting dust grains around the star. Although IRAS revealed that Fomalhaut is an infrared-excess source, these observations constitute the first detection of this extended, disk-like emission. Thus, after β Pic (also A3V), this is the second main-sequence star to have a detected disk, and the disk is the first one around a main-sequence star to be mapped in thermal emission. The fact that α PsA is 2.5 times closer to the earth than β Pic is makes it an ideal object for intensive study.”

SU TAURI This R-CrB variable is fading, as indicated by the following visual

magnitude estimates, communicated mainly by B. H. Granslo, Fjellhamar, Norway: Mar. 20.51 UT, 10.0 (M. Iida, Nagano, Japan); 23.42, 10.2 (M. Yamada, Ishikawa, Japan); 25.84, 10.4 (Granslo); 26.47, 10.5 (Yamada); 27.44, 10.8 (Yamada); 28.85, 11.3 (Granslo); 29.82, 11.6 (Granslo).

COMET SHOEMAKER-LEVY (1993e) A. Cochran, University of Texas, reports that she obtained spectra

(range 300-570 nm) of this comet with the large cassegrain spectrograph on the McDonald Observatory’s 2.7-m telescope on Mar. 28. The long slit was placed along the major axis of the object. A preliminary look at the raw data shows no obvious emissions. The comet was approximately uniformly bright over the inner 30′′. There was an approximately uniform section of lower intensity to a total length of 1′. A weak condensation was visible near the western end.

1993 March 30 Brian G. Marsden